OpenStudy (anonymous):

Calculate the average rate of change for the given graph from x = -2 to x = 0 and select the correct answer below

jimthompson5910 (jim_thompson5910):

Please attach the picture where it says "attach file" under the text box

jimthompson5910 (jim_thompson5910):

that works too

jimthompson5910 (jim_thompson5910):

`Calculate the average rate of change for the given graph from x = -2 to x = 0` that's the same as saying `calculate the slope between the two points (-2,-1) and (0,-1)`

OpenStudy (anonymous):

So what would I do?

jimthompson5910 (jim_thompson5910):

use this slope formula \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

jimthompson5910 (jim_thompson5910):

hopefully you've seen it before

OpenStudy (anonymous):

It would be -2?

jimthompson5910 (jim_thompson5910):

(x1,y1) = (-2,-1) (x2,y2) = (0,-1) \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] \[\Large m = \frac{-1-(-1)}{0-(-2)}\] \[\Large m = \frac{-1+1}{0+2}\] \[\Large m = \frac{0}{2}\] \[\Large m = 0\] The slope of the line through the two points (-2,-1) and (0,-1) is m = 0 So the average rate of change from x = -2 to x = 0 is 0

OpenStudy (anonymous):

I get it! Thanks!!!

jimthompson5910 (jim_thompson5910):

no problem